OpenStudy (anonymous):

The ratio of men to women working for a company 8 to 5. If there are 145 women working for the company, what is the total number of employees?

OpenStudy (solomonzelman):

\[\color{blue}{ \frac{8}{5}=\frac{x}{145} } \] solve for x. x is number of men. when you find x, \[\color{blue}{ x+145=total~~~number~~~of~~~workers } \]

OpenStudy (anonymous):

I got 232 o.O because I did 8*145/5

OpenStudy (solomonzelman):

232 for men?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

Yeah, correct. So the total number of workers, men and woman =?

OpenStudy (anonymous):

377

OpenStudy (solomonzelman):

Yup!

OpenStudy (anonymous):

I have another one similar to that and I'm having trouble with it

OpenStudy (solomonzelman):

Sure

OpenStudy (solomonzelman):

r u from Atlanta?

OpenStudy (anonymous):

It contains 3 liters of natural oil and 4 liters of synthetic oil and it has 378 Petrolyn oil, how many liters of synthetic oil are needed?

OpenStudy (solomonzelman):

needed for what?

OpenStudy (anonymous):

oops, it says in order to make 378 petrolyn oil how many liters of synthetic oil are needed

OpenStudy (solomonzelman):

I get what should be written, like what you meant to write, but I don't get the question?

OpenStudy (anonymous):

Petrolyn motor oil is a combination of natural oil and synthetic oil. It contains 3 liters of natural oil for every 4 liters of synthetic oil. In order to make 378 liters of Petrolyn oil, how many liters of synthetic oil are needed?

OpenStudy (solomonzelman):

Yes, that's much better!

OpenStudy (anonymous):

Haha :)

OpenStudy (solomonzelman):

Ok, so you have 3 natural + 4 synthetic = 7 Petrolyn 3n+4s=378

OpenStudy (solomonzelman):

I know it, but its really hard to explain.

OpenStudy (anonymous):

Is the answer 640?

OpenStudy (solomonzelman):

wait is it that 4 liters of synthetic and 3 of natural make 1 liter of petrolyn or 7 liters of petrolyn?

OpenStudy (anonymous):

I put 7 liters

OpenStudy (anonymous):

It's 7

OpenStudy (solomonzelman):

OK. so synthetic oil is going to 4/7 the ingredients.

OpenStudy (solomonzelman):

\[\color{blue}{ 378 \div 7 \times 4=216 } \]

OpenStudy (anonymous):

Thanks :) I will leave you alone now lol.